# Relative pose of three calibrated and partially calibrated cameras from four points using virtual correspondences

### v2

@@ 9.2 Noise experiments
We tested the performance of our solvers and the state-of-the-art algorithms w.r.t. increasing image noise. We used the SfM model of the botanical garden scene (randomly selected from all scenes) from the ETH3D dataset [53] to obtain instances of 5/6 points in three views by identifying images in the scene that share 3D points. Perfect noise-free correspondences are generated by projecting the 3D points into the images. We then add increasing amounts of normally distributed noise to these correspondences. We generated more than 9k instances, but show only 1k results per plot to avoid clutter. Note that the 4p3v(HC) solver was trained on the ETH3D dataset while our L-based solvers were trained on purely synthetic data.
@@
The results for increasing noise in the image points are shown in Figs. 6 and 7. The results are represented by boxplot function which shows the 25% to 75% quantiles as a box with a horizontal line at median. Crosses show data beyond 1.5 times the interquartile range. Let $e(\mathtt{R}_{ij})$ be the error of the estimated relative rotation between cameras $i$ and $j$, computed as the angle in the axis-angle representation of $\mathtt{R}_{ij}^{-1}\mathtt{R}_{ij}^{\mathtt{GT}}$ and let $e(\mathtt{t}_{ij})$ be the error of the estimated translation computed as the angle between the two unit vectors corresponding to the translations. Fig. 6 shows boxplots of pose errors measured in the same way as in our experiments in the main paper (cf. Sec. 4 in the main paper), i.e. as $\texttt{max}(\texttt{avg}(e(\mathtt{R}_{12}),e(\mathtt{R}_{13})),\texttt{avg}(e(\mathtt{t}_{12}),e(\mathtt{t}_{13})))$, for the calibrated 4p3v problem (left), and the partially calibrated 4p3vf problem (right). The errors are zoomed into an interesting interval and are shown as functions of varying noise from $0px$ to $4px$.
@@
Due to the approximate nature of the virtual correspondences, our newly proposed M-based and L-based solvers return non-zero errors for zero noise. However, at noise levels $\geq 1px$, our $\delta$-based solvers (both M and L), and for the calibrated case even the pure 4p3v(M) and 4p3v(L) solvers, return comparable or even better results than the 5pt+P3P and 6pt+P3P solvers. Note that the 5pt+P3P and 6pt+P3P solvers sample one/two more points (real correspondences) in the first two cameras, and these points are affected only by the considered noise. This shows that our predicted virtual correspondences are good approximations to real correspondences. For the calibrated case, the recent state-of-the-art solver [21] is failing in about 50% of the instances for noiseless data, even though the solver was trained on the ETH3D dataset. Thus, the median errors are significantly larger than the median errors of the remaining solvers.
@@
The rotation and translation errors in the first two views, i.e., $e(\mathtt{R}_{12})$ and $e(\mathtt{t}_{12})$, for both the calibrated (top row), and the partially calibrated case (bottom row) are shown in Fig. 7. For the partially calibrated case, our new solvers generate two approximate virtual correspondences in the first two views. Therefore, the 4p3vf(M) and 4p3vf(L) solvers have slightly larger errors than the 6pt+P3P solver for all considered noise levels. However, similarly to the pose errors in Figure 6, at noise levels $\geq 1px$ our $\delta$-based solvers (both M and L), and for the calibrated case even the pure 4p3v(M) and 4p3v(L) solvers, return comparable or even better results in the first two views than the 5pt [38] and 6pt [59] solvers, here represented by the results of 5pt+P3P and 6pt+P3P solvers. This shows an interesting potential of using our solvers for the two-view relative pose estimation problems by solving these problems from sub-minimal samples.
@@ 9.3 Outlier experiments
(a) 0.1 Inlier Ratio
(b) 0.2 Inlier Ratio
(c) 0.4 Inlier Ratio
Figure 8: Outlier Experiments: We tested the performance, in terms of the average percentage of inliers, of the solvers for different inlier ratios and for different numbers of RANSAC iterations. Results are shown for the Sacre Coeur scene from PhotoTourism dataset [54]. We plot the results for (a) 0.1 inlier ratio, (b) 0.2 inlier ratio, and (c) 0.4 inlier ratio.
(a) 0.1 Inlier Ratio
(b) 0.2 Inlier Ratio
(c) 0.4 Inlier Ratio
Figure 9: Outlier Experiments: We tested the performance, in terms of the average percentage of inliers, of the solvers for different inlier ratios and for different numbers of RANSAC iterations. Results are shown for the St Peters’ Square scene from PhotoTourism dataset [54]. We plot the results for (a) 0.1 inlier ratio, (b) 0.2 inlier ratio, and (c) 0.4 inlier ratio.
(a) 0.1 Inlier Ratio
(b) 0.2 Inlier Ratio
(c) 0.4 Inlier Ratio
Figure 10: Outlier Experiments: We tested the performance, in terms of the average pose error (maximum of rotation and translation errors), of the solvers for different inlier ratios and for different numbers of RANSAC iterations. Results are shown for the St. Peters’ Square scene from the PhotoTourism dataset [54]. We plot the results for (a) 0.1 inlier ratio, (b) 0.2 inlier ratio, and (c) 0.4 inlier ratio.
The outlier experiments are implemented as described in Section 5 the main paper. Here we provide additional plots for inlier ratios $\{0.1,0.2,0.4\}$. In Figures 8 and 9, we show the performance of each solver w.r.t. the percentage of inliers that they gather, on different RANSAC iterations, for the Sacre Coeur and St. Peters’ Square scenes, respectively. In Figure 10, we show the performance of the solvers w.r.t. average pose error on the St. Peters’ Square scene, similar to the one included in the main paper for Sacre Coeur. The pose error was measured in the same way as in our experiments in the main paper (cf. Sec. 4 in the main paper), i.e., as $\texttt{max}(\texttt{avg}(e(\mathtt{R}_{12}),e(\mathtt{R}_{13})),\texttt{avg}(e(\mathtt{t}_{12}),e(\mathtt{t}_{13})))$.
For low inlier ratios and in early RANSAC iterations, we observe that the $\delta$-based solvers perform the best, slightly better than 4p3v(M) and 4p3v(L). 5pt+P3P and 4p3v(HC) solvers show worse performance than our proposed solvers in this experiment. Increasing the inlier ratio up to 0.4, the performance of 5pt+P3P increases, becoming similar to the performance of 4p3v(M) and 4p3v(L). Even for an inlier ratio of 0.4, 4p3v(HC) exhibits a slower convergence due to the higher failure rate.
9.4 Timings

### v3

@@ 9.2 Noise experiments
We tested the performance of our solvers and the state-of-the-art algorithms w.r.t. increasing image noise. We used the SfM model of the botanical garden scene (randomly selected from all scenes) from the ETH3D dataset [50] to obtain instances of 5/6 points in three views by identifying images in the scene that share 3D points. Perfect noise-free correspondences are generated by projecting the 3D points into the images. We then add increasing amounts of normally distributed noise to these correspondences. We generated more than 9k instances, but show only 1k results per plot to avoid clutter. Note that the 4p3v(HC) solver was trained on the ETH3D dataset while our L-based solvers were trained on purely synthetic data. For the noise experiments we also test the joint 5PC solver, which operates on samples of 5 point correspondences, by estimating the essential matrices $\mathtt{E}_{12},\mathtt{E}_{13}$ independently, using the 5pt solver. Notice that due to estimating the two essential matrices independently of each other, the scales of both translations are indepent from each other. In contrast, the other solvers estimate the scale of the translation of the third camera relative to the scale of the translation between the first two cameras.
@@
The results for increasing noise in the image points are shown in Figs. 10 and 11. The results are represented by the boxplot function which shows the 25% to 75% quantiles as a box with a horizontal line at median. Crosses show data beyond 1.5 times the interquartile range. Let $\mathtt{R}_{err}^{ij}$ be the error of the estimated relative rotation between cameras $i$ and $j$, computed as the angle in the axis-angle representation of $\mathtt{R}_{ij}^{-1}\mathtt{R}_{ij}^{\mathtt{GT}}$ and let $\mathbf{t}_{err}^{ij}$ be the error of the estimated translation computed as the angle between the two unit vectors corresponding to the translations [20]. Fig. 10 shows boxplots of pose errors measured in the same way as in our experiments in the main paper (cf. Sec. 4 in the main paper), i.e., as $\text{max}\left(0.5(\mathtt{R}_{err}^{12}+\mathtt{R}_{err}^{13}),0.5(\mathbf{t}_{err}^{12}+\mathbf{t}_{err}^{13})\right)$, for the calibrated 4p3v problem (left), and the partially calibrated 4p3vf problem (right). The errors are zoomed into an interesting interval and are shown as functions of varying noise from $0px$ to $4px$.
@@
Due to the approximate nature of the virtual correspondences, our newly proposed M-based and L-based solvers exhibit non-zero errors for zero noise. However, at noise levels $\geq 1px$, our $\delta$-based solvers (both M and L), and for the calibrated case even the pure 4p3v(M) and 4p3v(L) solvers, return comparable or even better results than the 5pt+P3P and 6pt+P3P solvers. Note that the 5pt+P3P and 6pt+P3P solvers sample one/two more points (real correspondences) in the first two cameras, and these points are affected only by the considered noise. This shows that our predicted virtual correspondences are good approximations to real correspondences. For the calibrated case, the recent state-of-the-art solver [19] is failing in about 50% of the instances for noiseless data, even though the solver was trained on the ETH3D dataset. Thus, the median errors are significantly larger than the median errors of the remaining solvers.
@@
The rotation and translation errors in the first two views, i.e., $\mathtt{R}_{err}^{12}$ and $\mathbf{t}_{err}^{12}$, for both the calibrated (top row), and the partially calibrated case (bottom row) are shown in Fig. 11. For the partially calibrated case, our new solvers generate two approximate virtual correspondences in the first two views. Therefore, the 4p3vf(M) and 4p3vf(L) solvers have slightly larger errors than the 6pt+P3P solver for all considered noise levels. However, similarly to the pose errors in Figure 10, at noise levels $\geq 2px$ our $\delta$-based solvers (both M and L) return comparable or even better results in the first two views than the 5pt [34] and 6pt [56] solvers, here represented by the results of 5pt+P3P and 6pt+P3P solvers. For the calibrated case even the pure 4p3v(M) and 4p3v(L) solvers, without the offset $\delta$, perform comparably well as the 5pt solver. This shows an interesting potential of using our solvers for the two-view relative pose estimation problems by solving these problems from sub-minimal samples.
(a) Phototourism [20]
(b) Cambridge Landmarks [21]
Figure 12: Speed-accuracy trade-off on (a) all scenes from Phototourism [20] except St. Peter’s Square, (b) 5 Cambridge Landmarks [21] scenes. We report the AUC@10 using the alternative definition of pose error (3). We vary the number of Poselib RANSAC iterations (100, 200, 500, 1000, 2000, 5000, 10000). Runtimes are averaged over all image triplets. Legend is the same as in Fig. 14.
Phototourism [20]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (s) $\downarrow$
4p3v(HC) [19] 5.23 1.89 43.36 62.83 76.97 2.95
5pt+P3P 5.00 1.85 43.99 63.35 77.39 2.78
4p3v(M) 5.11 1.94 43.03 62.87 77.15 2.23
4p3v(M)+R 5.07 1.91 43.30 63.14 77.32 2.47
4p3v(M)+R+F 5.05 1.89 43.41 63.24 77.39 2.42
4p3v(M±$\pm\delta$) 5.02 1.92 43.24 63.15 77.44 2.25
4p3v(M±$\pm\delta$)+R 4.96 1.90 43.57 63.48 77.67 2.53
4p3v(M±$\pm\delta$)+R+F 5.00 1.89 43.51 63.38 77.56 2.41
4p3v(L) 5.46 1.93 42.82 62.07 76.25 2.88
4p3v(L)+R 5.05 1.91 43.28 63.19 77.42 2.50
4p3v(O)+R 4.70 1.81 44.73 64.50 78.44 2.38
4p3v(NO)+R 4.24 1.74 46.01 65.97 79.74 2.39
Table 7: Results for different methods implemented in the GC-RANSAC framework [2] for all scenes from the PhotoTourism [20] dataset. We mark the best and second best results (excluding oracle solvers).
Phototourism [20]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (ms) $\downarrow$
4p3v(HC) [19] 11.41 3.89 37.90 53.76 68.08 176.45
5pt+P3P 19.88 3.41 41.36 57.36 71.13 105.50
4p3v(M) 11.65 4.21 35.38 51.73 66.77 176.77
4p3v(M)+R 10.52 3.47 41.01 56.62 70.25 174.94
4p3v(M)+R+F 10.79 3.58 40.09 55.73 69.56 130.54
4p3v(M±$\pm\delta$) 10.44 3.72 38.82 55.07 69.51 172.06
4p3v(M±$\pm\delta$)+R 19.84 3.27 42.62 58.30 71.74 189.21
4p3v(M±$\pm\delta$)+R+F 10.14 3.40 41.45 57.24 70.92 175.78
4p3v(L) 11.56 4.21 35.32 51.77 66.84 376.31
4p3v(L)+R 10.48 3.47 41.02 56.60 70.25 297.88
4p3v(L±$\pm\delta$init)+R 19.86 3.28 42.60 58.26 71.71 730.90
4p3v(O) 19.64 3.29 42.60 58.54 72.09 158.22
4p3v(O)+R 19.50 3.13 44.01 59.67 72.81 186.30
4p3v(O)+R+F 19.45 3.16 43.82 59.55 72.76 136.36
Cambridge Landmarks [21]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (ms) $\downarrow$
4p3v(HC) [19] 15.12 5.51 24.50 43.42 61.10 164.49
5pt+P3P 13.38 5.17 25.94 45.38 63.16 148.53
4p3v(M) 15.31 5.68 22.93 42.06 60.26 138.33
4p3v(M)+R 14.28 5.29 25.37 44.57 62.33 134.84
4p3v(M)+R+F 14.60 5.39 24.89 43.97 61.71 116.51
4p3v(M±$\pm\delta$) 14.18 5.36 24.72 44.12 62.15 181.21
4p3v(M±$\pm\delta$)+R 13.67 5.11 26.17 45.56 63.33 184.98
4p3v(M±$\pm\delta$)+R+F 13.89 5.23 25.43 44.78 62.70 138.45
4p3v(L) 15.24 5.72 22.98 42.15 60.35 232.13
4p3v(L)+R 14.23 5.27 25.33 44.49 62.30 177.06
4p3v(L±$\pm\delta$init)+R 13.63 5.10 26.11 45.58 63.35 395.43
4p3v(O) 13.85 5.22 25.61 45.19 63.11 130.43
4p3v(O)+R 13.51 5.07 26.60 46.05 63.76 134.68
4p3v(O)+R+F 13.47 5.10 26.47 45.98 63.71 116.41
Table 8: Results for different solvers implemented in the PoseLib framework [26] on all scenes from the PhotoTourism [20] and 5 scenes from the Cambridge Landmarks [21] datasets with the alternative definition of pose error (3). We mark the best and second best results (excluding oracle solvers). Reported runtimes are for the whole RANSAC.
9.3 Alternative evaluation measure
For the evaluation in the main paper, we defined the pose error as $\text{max}\left(0.5(\mathtt{R}_{err}^{12}+\mathtt{R}_{err}^{13}),0.5(\mathbf{t}_{err}^{12}+\mathbf{t}_{err}^{13})\right)$, where $\mathtt{R}_{err}^{ij}$ and $\mathbf{t}_{err}^{ij}$ are the angular errors of rotation and translation for camera pair $ij$ in degrees. The 4p3v problem also includes the estimation of $\mathtt{R}_{23}$ and $\mathbf{t}_{23}$ since the relative scale of $\mathbf{t}_{12}$ and $\mathbf{t}_{13}$ is recovered. We therefore also presents results for the pose error defined as
$P_{err}=\text{max}\left(\mathtt{R}_{err}^{12},\mathtt{R}_{err}^{13},\mathtt{R}_{err}^{23},\mathbf{t}_{err}^{12},\mathbf{t}_{err}^{13},\mathbf{t}_{err}^{23}\right)\enspace.$ (3)
The results equivalent to Tab. 2 from the main paper using this pose error definition are presented in Tab. 6. A speed-accuracy comparison equivalent to Fig. 3 in the main is presented in Fig. 12. The overall comparison of the methods remains the same under both the metric used in the main paper and the alternative described in this section.
9.4 GC-RANSAC
We evaluated and compared our proposed solvers against the state-of-the-art solver, also in the GC-RANSAC [2] framework. Results for all scenes of the PhotoTourism dataset [20] are presented in Tab. 7.
4p3v(M)4p3v(M)+R4p3v(M)+R+F4p3v(M±$\pm\delta$)4p3v(M±$\pm\delta$)+R4p3v(M±$\pm\delta$)+R+F4p3v(HC)5pt+P3P4p3v(O)+R+F
St. Mary’s Church [21]
(a) Original Matches
(b) Synth - 0.6 inlier ratio
(c) Synth - 0.4 inlier ratio
(d) Synth - 0.2 inlier ratio
Sacre Coeur [20]
(e) Original Matches
(f) Synth - 0.6 inlier ratio
(g) Synth - 0.4 inlier ratio
(h) Synth - 0.2 inlier ratio
Figure 13: Outlier experiments for St. Mary’s Church scene from Cambridge Landmarks [21] (a-d) and Sacre Coeur scene from Photorouism [20] (e-h). We show the speed-accuracy trade-off evaluation (see Fig. 14) for the original matches (a,e) and for synthetic scenario where we kept the matches that are inliers w.r.t. ground truth poses and added randomly distributed noisy matches to obtain desired inlier ratio (b-d,f-h).
@@ 9.5 Outlier experiments
Fig. 4 in the main paper presented results obtained by synthetically removing outlier matches based on ground truth pose information and replacing them with outliers distributed uniformly at random to reach a given inlier ratio for all image triplets. Here we provide additional plots for two scenes: St Mary’s Church from Cambridge Landmarks [21] and Sacre Coeur from Phototourism [20]. The results for the speed-accuracy evaluation are shown in Fig. 13. Consistent with the results in the main paper, these graphs show the potential for our M-based solvers to perform better than the the 5pt+P3P baseline in scenarios with low inlier ratios.
4p3v(M)4p3v(M)+R4p3v(M)+R+F4p3v(M±$\pm\delta$)4p3v(M±$\pm\delta$)+R4p3v(M±$\pm\delta$)+R+F4p3v(HC)5pt+P3P4p3v(O)+R+F
(a) Great Court
(b) Shop Facade
(c) Old Hospital
(d) Buckingham Palace
(e) Colosseum Exterior
(f) Grand Place Brussels
(g) Notre Dame
(h) Palace of Westminster
(i) Pantheon Exterior
(j) Taj Mahal
(k) Temple Nara
(l) Trevi Fountain
Figure 14: Results for individual scenes from the Cambrdige Landmarks [21] (a-c) and Phototourism [20] (d-l) datasets which were not presented in the main paper or Fig. 13. We report the AUC@10 of the pose error and vary the number of Poselib RANSAC iterations (100, 200, 500, 1000, 2000, 5000, 10000). Runtimes are averaged over all image triplets.
9.6 Detailed experiments on Real Data
Results for individual scenes.  Fig. 3 in the main paper shows results on all PhotoTourism scenes (except St. Peter’s Square), the 5 Cambridge Landmarks scenes we consider, and one individual scene from each dataset. In Fig. 14, we provide results for the accuracy-speed trade-off evaluation for more evaluation scenes on both PhotoTourism [20] and Cambridge Landmarks [21]. As discussed and shown in the main paper, the performance of our M-based and L-based solvers is scene-dependent. This can also be seen in Fig. 14, where for some scenes, the 4p3v(M)+R and 4p3v(M±$\pm\delta$)+R solvers perform worse than the 5pt+P3P solver (Shop Facade, Palace of Westminster, Trevi Fountain). However, for the majority of the scenes, our solvers perform similar to 5pt+P3P (Old Hospital) or even outperform 5pt+P3P (Great Court, Buckingham Palace, Colosseum Exterior, Grand Place Brussels, Notre Dame, Pantheon Exterior, Taj Mahal, Temple Nara). Overall, the results validate the practical viability of our solvers in a time-constrained setting.
Results for L-based solvers.  In Tab. 9 we provide the same results as shown in Tab. 3 in the main paper, including also more variants for L-based solvers, i.e. 4p3v(L)+R+F, 4p3v(L±$\pm\delta$), 4p3v(L±$\pm\delta$)+R, 4p3v(L±$\pm\delta$)+R+F, 4p3v(L±$\pm\delta$init), 4p3v(L±$\pm\delta$init)+R, and 4p3v(L±$\pm\delta$init)+R+F.
Phototourism [20]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (ms) $\downarrow$
4p3v(HC) [19] 7.17 2.34 52.74 66.63 77.86 176.45
5pt+P3P 5.99 2.00 57.31 70.54 80.81 105.50
4p3v(M) 7.17 2.49 50.96 65.46 77.32 176.77
4p3v(M)+R 6.39 2.00 56.92 69.90 80.17 174.94
4p3v(M)+R+F 6.59 2.07 55.90 69.06 79.56 130.54
4p3v(M±$\pm\delta$) 6.39 2.19 54.70 68.58 79.59 172.06
4p3v(M±$\pm\delta$)+R 5.97 1.89 58.65 71.43 81.35 189.21
4p3v(M±$\pm\delta$)+R+F 6.15 1.97 57.42 70.47 80.69 175.78
4p3v(L) 7.12 2.50 51.00 65.57 77.42 376.31
4p3v(L)+R 6.35 2.00 56.88 69.88 80.15 297.88
4p3v(L)+R+F 6.55 2.07 55.88 69.08 79.55 255.93
4p3v(L±$\pm\delta$) 6.39 2.19 54.70 68.58 79.59 172.05
4p3v(L±$\pm\delta$)+R 5.97 1.89 58.65 71.43 81.35 189.28
4p3v(L±$\pm\delta$)+R+F 6.15 1.97 57.42 70.47 80.69 175.76
4p3v(L±$\pm\delta$init) 5.97 1.91 58.36 71.25 81.26 679.80
4p3v(L±$\pm\delta$init)+R 5.97 1.89 58.63 71.41 81.34 730.90
4p3v(L±$\pm\delta$init)+R+F 6.04 1.92 58.11 70.97 81.02 605.86
4p3v(O) 5.82 1.90 58.91 71.84 81.70 158.22
4p3v(O)+R 5.73 1.80 60.23 72.75 82.21 186.30
4p3v(O)+R+F 5.72 1.82 59.97 72.58 82.13 136.36
Cambridge Landmarks [21]
Estimator AVG $(^{\circ})$downarrow MED $(^{\circ})$downarrow AUC@5 $\uparrow$ @10 $\uparrow$ @20 $\uparrow$ Runtime (ms) $\downarrow$
4p3v(HC) [19] 9.69 3.31 40.96 58.84 72.83 164.49
5pt+P3P 8.16 3.05 43.79 61.61 75.30 148.53
4p3v(M) 9.61 3.42 39.71 58.08 72.54 138.33
4p3v(M)+R 8.77 3.11 42.98 60.90 74.59 134.84
4p3v(M)+R+F 9.03 3.17 42.31 60.18 74.03 116.51
4p3v(M±$\pm\delta$) 8.75 3.21 42.11 60.37 74.38 181.21
4p3v(M±$\pm\delta$)+R 8.32 3.01 44.17 62.04 75.60 184.98
4p3v(M±$\pm\delta$)+R+F 8.47 3.08 43.21 61.22 75.03 138.45
4p3v(L) 9.58 3.44 39.79 58.17 72.62 232.13
4p3v(L)+R 8.75 3.09 42.94 60.86 74.60 177.06
4p3v(L)+R+F 8.92 3.17 42.28 60.19 74.08 163.95
4p3v(L±$\pm\delta$) 8.75 3.21 42.11 60.37 74.38 181.20
4p3v(L±$\pm\delta$)+R 8.32 3.01 44.17 62.04 75.60 184.89
4p3v(L±$\pm\delta$)+R+F 8.47 3.08 43.21 61.22 75.03 138.49
4p3v(L±$\pm\delta$init) 8.43 3.05 43.80 61.70 75.34 379.49
4p3v(L±$\pm\delta$init)+R 8.28 3.00 44.26 62.10 75.62 395.43
4p3v(L±$\pm\delta$init)+R+F 8.33 3.04 43.82 61.71 75.35 349.45
4p3v(O) 8.62 3.07 43.58 61.65 75.30 130.43
4p3v(O)+R 8.39 2.94 44.80 62.54 75.93 134.68
4p3v(O)+R+F 8.36 2.97 44.64 62.38 75.82 116.41
Table 9: Results for different solvers implemented in the PoseLib framework [26] on all scenes from the PhotoTourism [20] and 5 scenes from the Cambridge Landmarks [21] datasets including all evaluated variants of the proposed solvers. We mark the best and second best results (excluding oracle solvers). Reported runtimes are for the whole RANSAC.
5pt+P3P 4p3v(HC) 4p3v(M) 4p3v(M±$\pm\delta$) 4p3v(L) 4p3v(L±$\pm\delta$) 4p3v(L±$\pm\delta$init)
Time ($\mu$s) 77.90 66.06 83.92 218.71 450.26 511.28 1130.31
Table 10: The average run-time, averaged over more than 10k instances of the Sacre Coeur scene of the PhotoTourism dataset [51], of the solvers for the calibrated case.